OpenStudy (anonymous):

9 3/4 percent converted to a fraction equals

OpenStudy (anonymous):

write 9 as 36 / 4 and add to 3/4 to get an improper fraction

OpenStudy (anonymous):

\[9\text{ and }\frac{3}{4} = 9 + \frac{3}{4} = \frac{36}{4} + \frac{3}{4} = ?\]